# references.bib

### v3

@@ 4.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-GE
A possible argument to achieve more accurate ML models is to build models only using the data source that one has control over, which would be the GE Data in the current study. To demonstrate the advantages of using multi-source data fusion through LVGP over a single-source model for the specified source of interest, the prediction performances of LVGP and the GP model trained on the GE Data only, which was previously denoted as GP-GE, are compared. Table 4 (Model Types: LVGP and GP-GE) and Panels B and C of Figure 6 demonstrate the prediction metrics and the response surface obtained from the models. Due to the differences in the data used to train models (multi-source vs single-source), testing accuracy of the models are studied on the common 36 GE Data testing set. The NMRSE prediction metrics from Table 4 demonstrate that incorporating other available sources into the modeling yield very similar predictive performances, with slightly better results towards GP-GE model. However, comparing the 45-degree parity plots between the two models (first row of Figure 6), we observe that the prediction uncertainties provided by the LVGP are much smaller compared to the prediction uncertainties provided by the GP-GE model. A possible explanation for this phenomenon could be that the knowledge incorporated from other sources contributes to the learning of the material system and therefore increases the confidence in the model when making predictions. Furthermore, the response surfaces between the two models show high similarity, indicating that the LVGP was able to capture the response surface of the specific source even when other data sources exist. The final very important observation is that the predictions provided by the LVGP model possess larger and lower maximum mean prediction values with lower uncertainties (standard deviation predictions) at different locations compared to both GP and GP-GE models. This observation demonstrates that the LVGP is able can perform extrapolation with much higher confidence by learning from other sources. This result could be significant and advantageous for materials design applications as it can give a multi-step head start to the design optimization task through more confident and accurate predictions.
@@ 4.4.3 Interpretability from the Latent Variables
A further validation study is done by separating the GE Data into two different sources as GE Data 1 and GE Data 2 and retraining the LVGP model to observe the latent space again. The aim of this validation study is to observe the new latent space when the same underlying experimental procedures are implemented for two resources along with other sources that have different unknown underlying procedures. The new constructed latent space is shown in Figure 8b. A significant observation from the validation study is that the latent variables of GE Data 1 and GE Data 2 are located adjacent to each other. This result was expected since the same experimental procedures are implemented for both sources. However, the results also once again demonstrated and validated that the LVGP model is able to capture the differences in the known and unknown underlying parameters that influence the material system and learn the relationships between the sources. This remarkable capability provided by the LVGP can be highly important for further downstream tasks including but not limited to model validation, dimension reduction, sensitivity analysis, and more. It is important to highlight that without using the LVGP approach proposed in this paper, it would be difficult to learn and assess the relationships between the data sources.
@@ 4.4.4 Enhancing Modeling Capabilities Through Targeted Source Selection Based On Latent Variables
So far, it has been demonstrated that the LVGP modeling brings numerous advantages to multi-source data fusion compared to some of the alternative approaches. One of the key advantages of the LVGP comes from the physically interpretable latent spaces that can possibly explain the relationships and differences between different sources of information. As a result, it is proposed that the latent space provided by the LVGP model can act as both a validation and an anomaly detection mechanism to remove the data samples from the sources that could potentially harm the ML model that will be used for further optimization tasks on the source of information that one has control over. For the data sources available in the FeCrAl alloy study, the authors generated and only have further control over the GE Data source. Therefore, the further material design optimization tasks can only take place using the experimental procedures available at GE Research. However, to improve the quality of the ML modelling, numerous data sources have also been collected and used for LVGP model building. Looking at the latent spaces from LVGP model in Figure 8a, it is observed that some data sources, Yang, Kobayashi, and Pinkas to be specific, are located at a much further point (higher D values) compared to the other sources when GE Data is selected as the reference point. The stark differences are also observed in the source-specific response surfaces learned and provided by the LVGP (Figure 7). As discussed before, this result can indicate that the experimental procedures taken place by those sources can be different than the ones that are implemented for GE Data. However, this does not necessarily state that there are fundamental errors or data quality issues regarding the three data sources but could potentially indicate that the sources closer to the GE Data can contribute to model accuracy whereas others can harm the ML model for predictions on GE Data. Therefore, to investigate the influence of targeted source selection on the separate testing GE Data set, the three sources that are further from the GE Data (Yang, Kobayashi, and Pinkas), is removed from the training sets and the LVGP model is re-trained. For ease of notation, the new LVGP model is denoted as LVGP-Target. Table 4 shows the comparison between the prediction accuracy made by the three models, LVGP, GP-GE, and LVGP-Target on the separate 36 GE Data testing set. Although both LVGP and GP-GE have already demonstrated remarkable accuracy results on the GE Data, LVGP-Target provides notable improvement in accuracy, specifically reducing NRMSE by almost 20% and 7% in comparison to LVGP and GP-GE models, respectively. The parity plot of the predictions made by the LVGP-Target model are shown in Supplementary Figure 2. Considering that the evaluation of properties through experimentation can be relatively expensive, the notable increase in accuracy provided by the targeted source selection through LVGP can provide significant cost and time savings for further downstream tasks. Overall, in addition to providing accurate data fusion modeling, LVGP can be employed as a validation and data filtering mechanism to improve multi-source modeling.
@@ 5.3 Model Development
Similar to the FeCrAl model development approach, GP and LVGP models are trained on the collected data. The common inputs for the GP and LVGP models are the composition of the alloy constituents (Samarium, Zirconium, Copper, Iron, Cerium, Cobalt) defined as the weight percentages (wt%) and the testing temperature of the alloy, leading to seven (7) unique input parameters. The energy density and power capacity of the magnet are experimentally measured and quantified through energy product (EP). As with the previous case study, the inputs for the LVGP contain an additional qualitative input characterizing each source of data as distinct categorical level along all the other 7 inputs used for the GP model. The outputs for both the models remain the same.
@@ 5.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-Prop
To overcome the challenge of performing optimization in the unknown information source space, one can argue to build GP models on each information source individually. To demonstrate the effectiveness of multi-source LVGP once again, it is compared against a single-source GP model that is built on the Prop SmCeCoFeCu source only. The single-source GP model, which is previously denoted as GP-Prop, uses the entire available training data from the Prop SmCeCoFeCu source (24 samples) for model building, and later tested on the separate 6 samples obtained from the Prop SmCeCoFeCu once again. The prediction metrics and the parity plots on testing data are shown in Table 6 (Model Types: LVGP and GP-Prop) and Panels B and C of Figure 9, respectively. From the NRMSE and parity plot results, a clear advantage of using LVGP is observed as it provides much better predictions with lower prediction uncertainties. Furthermore, looking at the response surfaces of both models, Panels B and C of Figure 9, a tone stark difference is seen in the uncertainty. For LVGP, the uncertainty for majority of the surface is at 0.5 MGOe. For th GP-Prop, the uncertainity for the majority of the surface is at 1.5 MGOe, three times of what is observed for LVGP. The high reduction in uncertainty of LVGP model comes from adding in other data-sources. This phenomenon could be a crucial differentiator when designing next generation materials with restrictions on data-availability driven by manufacturing and experimental testing capabilities.
@@ 6 Conclusion
AI and ML frameworks have had a strong impact in the science and engineering communities. On the other hand, with the ever-rising application of data-driven surrogates build on multiple sources of data that it is gathered either experimentally or computationally, not much attention has been paid to the quality, completeness, and comprehensiveness of the excluded and unknown underlying parameters involved in the data generation process. These parameters vary from differences in manufacturing process, testing technique, and more. Through the research done in this paper, an interpretable multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed and extensively studied. Through LVGP models, a categorical variable was introduced to represent sources of information during model training and predictions. A low-dimensional latent space is then learned to incorporate the categorical source variable into multi-source data fusion. Through latent variable representation, both the excluded and underlying unknown physical parameter variations among the different sources are captured and introduced into the modelling, making the multi-source ML model interpretable and aware of the differences and similarities between the information sources. The inclusion of the data sources not only leads to more accurate models but also enables the study of source specific response surfaces, material optimization, and sensitivity analysis. Compared to using single source data for ML, the proposed data fusion method using multi-source information can provide a better prediction for sparse-data problems by taking advantage of the information of correlations and relationships among different data sources. With the quantified uncertainty from LVGP for the specified information source, the approach can further guide adaptive sampling or active learning when allocating resources for data collection. Additionally, a novel dissimilarity metric is also defined through the latent variables of the LVGP models that can help understand the aggregated differences and similarities in the information sources. The proposed approach is applied on and analyzed through two mathematical (representative parabola and 2D Ackley function) and two materials design (design of FeCrAl alloys and design of SmCoFe alloys) case studies. Finally, a novel application of the framework would involve performing materials design and development through autonomous experimentation studies where multiple sources of experimental data are gathered to guide the study. It is envisioned that the proposed multi-source framework is not limited to experimental or computational information sources and can be further extended to mixture of experimental and computational sources in numerous engineering and science applications.
@@ Acknowledgments
Figure 1: Cross Validation performance on FeCrAl Alloy dataset using (a) GP and (b) LVGP models. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainity associated with each prediction
Figure 2: Training and testing performance on FeCrAl Alloy dataset using the LVGP-Target model. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainity associated with each prediction

### v4

@@ 4.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-GE
A possible argument to achieve more accurate ML models is to build models only using the data source that one has control over, which would be the GE Data in the current study. To demonstrate the advantages of using multi-source data fusion through LVGP over a single-source model for the specified source of interest, the prediction performances of LVGP and the GP model trained on the GE Data only, which was previously denoted as GP-GE, are compared. Table 4 (Model Types: LVGP and GP-GE) and Panels B and C of Figure 6 demonstrate the prediction metrics and the response surface obtained from the models. Due to the differences in the data used to train models (multi-source vs single-source), testing accuracy of the models are studied on the common 36 GE Data testing set. The NMRSE prediction metrics from Table 4 demonstrate that incorporating other available sources into the modeling yield very similar predictive performances, with slightly better results towards GP-GE model. However, comparing the 45-degree parity plots between the two models (first row of Figure 6), we observe that the prediction uncertainties provided by the LVGP are much smaller compared to the prediction uncertainties provided by the GP-GE model. The distribution of predicted uncertainties for both models are provided in Supplementary Figure 2. A possible explanation for this phenomenon could be that the knowledge incorporated from other sources contributes to the learning of the material system and therefore increases the confidence in the model when making predictions. Furthermore, the response surfaces between the two models show high similarity, indicating that the LVGP was able to capture the response surface of the specific source even when other data sources exist. The final very important observation is that the predictions provided by the LVGP model possess larger and lower maximum mean prediction values with lower uncertainties (standard deviation predictions) at different locations compared to both GP and GP-GE models. This observation demonstrates that the LVGP is able can perform extrapolation with much higher confidence by learning from other sources. This result could be significant and advantageous for materials design applications as it can give a multi-step head start to the design optimization task through more confident and accurate predictions.
@@ 4.4.3 Interpretability from the Latent Variables
A further validation study is done by splitting the GE Data into two distinct sources, assigning them different categorical levels as GE Data 1 and GE Data 2, and retraining the LVGP model to observe the latent space again. The aim of this validation study is to observe the new latent space when the same underlying experimental procedures are implemented for two resources along with other sources that have different unknown underlying procedures. The new constructed latent space is shown in Figure 8b. A significant observation from the validation study is that the latent variables of GE Data 1 and GE Data 2 are located adjacent to each other. This result was expected since the same experimental procedures are implemented for both sources. However, the results also once again demonstrated and validated that the LVGP model is able to capture the differences in the known and unknown underlying parameters that influence the material system and learn the relationships between the sources. This remarkable capability provided by the LVGP can be highly important for further downstream tasks including but not limited to model validation, dimension reduction, sensitivity analysis, and more. It is important to highlight that without using the LVGP approach proposed in this paper, it would be difficult to learn and assess the relationships between the data sources.
@@ 4.4.4 Enhancing Modeling Capabilities Through Targeted Source Selection Based On Latent Variables
So far, it has been demonstrated that the LVGP modeling brings numerous advantages to multi-source data fusion compared to some of the alternative approaches. One of the key advantages of the LVGP comes from the physically interpretable latent spaces that can possibly explain the relationships and differences between different sources of information. As a result, it is proposed that the latent space provided by the LVGP model can act as both a validation and an anomaly detection mechanism to remove the data samples from the sources that could potentially harm the ML model that will be used for further optimization tasks on the source of information that one has control over. For the data sources available in the FeCrAl alloy study, the authors generated and only have further control over the GE Data source. Therefore, the further material design optimization tasks can only take place using the experimental procedures available at GE Research. However, to improve the quality of the ML modelling, numerous data sources have also been collected and used for LVGP model building. Looking at the latent spaces from LVGP model in Figure 8a, it is observed that some data sources, Yang, Kobayashi, and Pinkas to be specific, are located at a much further point (higher D values) compared to the other sources when GE Data is selected as the reference point. The stark differences are also observed in the source-specific response surfaces learned and provided by the LVGP (Figure 7). As discussed before, this result can indicate that the experimental procedures taken place by those sources can be different than the ones that are implemented for GE Data. However, this does not necessarily state that there are fundamental errors or data quality issues regarding the three data sources but could potentially indicate that the sources closer to the GE Data can contribute to model accuracy whereas others can harm the ML model for predictions on GE Data. Therefore, to investigate the influence of targeted source selection on the separate GE Data testing set, the three sources that are further from the GE Data (Yang, Kobayashi, and Pinkas), are removed from the training sets and the LVGP model is re-trained with the remaining 229 samples and evaluated on the same GE Data testing set. For ease of notation, the new LVGP model is denoted as LVGP-Target. Table 4 shows the comparison between the prediction accuracy made by the three models, LVGP, GP-GE, and LVGP-Target on the separate 36 GE Data testing set. Although both LVGP and GP-GE have already demonstrated remarkable accuracy results on the GE Data, LVGP-Target provides notable improvement in accuracy, specifically reducing NRMSE by almost 20% and 7% in comparison to LVGP and GP-GE models, respectively. The parity plot of the predictions made by the LVGP-Target model are shown in Supplementary Figure 3. Considering that the evaluation of properties through experimentation can be relatively expensive, the notable increase in accuracy provided by the targeted source selection through LVGP can provide significant cost and time savings for further downstream tasks. Overall, in addition to providing accurate data fusion modeling, LVGP can be employed as a validation and data filtering mechanism to improve multi-source modeling.
@@ 5.3 Model Development
Similar to the FeCrAl model development approach, GP and LVGP models are trained on the collected data. The common inputs for the GP and LVGP models are the composition of the alloy constituents (Samarium, Zirconium, Copper, Iron, Cerium, Cobalt) defined as the weight percentages (wt%) and the testing temperature of the alloy, leading to seven (7) unique input parameters. The property of interest (output) for this case study is the energy product (EP) which is an experimentally measured metric that reflects the energy density and power capacity of the magnet. As with the previous case study, the inputs for the LVGP contain an additional qualitative input characterizing each source of data as distinct categorical level along all the other 7 inputs used for the GP model. The outputs for both the models remain the same.
@@ 5.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-Prop
To overcome the challenge of performing optimization in the unknown information source space, one can argue to build GP models on each information source individually. To demonstrate the effectiveness of multi-source LVGP once again, it is compared against a single-source GP model that is built on the Prop SmCeCoFeCu source only. The single-source GP model, which is previously denoted as GP-Prop, uses the entire available training data from the Prop SmCeCoFeCu source (24 samples) for model building, and later tested on the separate 6 samples obtained from the Prop SmCeCoFeCu once again. The prediction metrics and the parity plots on testing data are shown in Table 6 (Model Types: LVGP and GP-Prop) and Panels B and C of Figure 9, respectively. From the NRMSE metric, parity plot, and predicted uncertainty distributions (Supplementary Figure 5), the clear advantage of using LVGP is observed as it provides much better predictions with lower prediction uncertainties. Furthermore, looking at the response surfaces of both models, Panels B and C of Figure 9, a tone stark difference is seen in the uncertainty. For LVGP, the uncertainty for majority of the surface is at 0.5 MGOe. For the GP-Prop, the uncertainty for the majority of the surface is at 1.5 MGOe, three times of what is observed for LVGP. The high reduction in uncertainty of LVGP model comes from adding in other data sources. This phenomenon could be a crucial differentiator when designing next generation materials with restrictions on data-availability driven by manufacturing and experimental testing capabilities.
@@ 6 Conclusion
AI and ML frameworks have had a strong impact in the science and engineering communities. On the other hand, with the ever-rising application of data-driven surrogates build on multiple sources of data that it is gathered either experimentally or computationally, not much attention has been paid to the quality, completeness, and comprehensiveness of the excluded and unknown underlying parameters involved in the data generation process. These parameters vary from differences in manufacturing process, testing technique, and more. Through the research done in this paper, an interpretable multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed and extensively studied. Through LVGP models, a categorical variable was introduced to represent sources of information during model training and predictions. A low-dimensional latent space is then learned to incorporate the categorical source variable into multi-source data fusion. Through latent variable representation, both the excluded and underlying unknown physical parameter variations among the different sources are captured and introduced into the modelling, making the multi-source ML model interpretable and aware of the differences and similarities between the information sources. The inclusion of the data sources not only leads to more accurate models but also enables the study of source specific response surfaces, material optimization, and sensitivity analysis. Compared to using single source data for ML, the proposed data fusion method using multi-source information can provide a better prediction for sparse-data problems by taking advantage of the information of correlations and relationships among different data sources. Additionally, a novel dissimilarity metric is also defined through the latent variables of the LVGP models that can help understand the aggregated differences and similarities in the information sources. The proposed approach is applied on and analyzed through two mathematical (representative parabola and 2D Ackley function) and two materials design (design of FeCrAl alloys and design of SmCoFe alloys) case studies.
Although the demonstrated method is focused on the material science domain, the proposed approach of leveraging LVGP for data fusion modeling can be applicable to other several domains of engineering where multiple source scenarios are observed frequently. The definition of a source, set of data points or a data generating process with different mapping behavior between the fixed set of inputs and outputs, is what permeates the approach to domains. This definition of a source become inclusive of scenarios where sources are differentiated by fidelity to evaluate a design (e.g. potential flow solvers, Reynolds-Averaged Navier-Stokes and Large Eddy Simulation, for computational fluid dynamics applications, static and dynamic analysis for structural design), by differences in operational history of manufacturing systems (e.g., milling machine, binder jet machine) leading to machine to machine variations, and by differences in health of mechanical system (e.g., jet engines, gear systems) leading to performance variations. For these applications, the nuances between the sources does not need to be quantified directly and can be simply tagged as categorical variable into the LVGP model where the latent variables of the LVGP can internally learn the differences in sources. With the quantified uncertainty from LVGP for the specified information source, the approach can further guide adaptive sampling or active learning when allocating resources for data collection. As such, a novel application of the framework would involve performing materials design and development through autonomous experimentation studies where multiple sources of experimental data are gathered to guide the study. It is envisioned that the proposed multi-source framework is not limited to experimental or computational information sources and can be further extended to mixture of experimental and computational sources in numerous engineering and science applications.
Finally, a key challenge in the presented method lies when there are high number of sources present, potentially more than 15 sources. In that case, the LVGP model might struggle to converge to a final latent space due to high number of parameter (latent variables) estimations during model fitting. However, it was previously shown that even when a large number of categorical variable options are present, further design optimization campaign works efficiently to identify top designs for metal-organic framework materials [comlek2023rapid] . Due to time limitations for training GP models with the presence of large datasets, a future research direction for multi-source modeling through LVGP involves modeling through sparse-variational LVGP [wang2022scalable]. Furthermore, as each information source could also contain different input parametrizations or spaces (heterogeneous inputs), extending LVGP to handle heterogeneous cases can benefit the numerous engineering applications [comlek2024heterogenous].
Data Availability
The MATLAB and python code for LVGP used in this study can be found at https://github.com/ideal-nu/MOF-LVGP-MOBBO [comlek2023rapid] and https://github.com/syerramilli/lvgp-bayes [yerramilli2023fully].
@@ Acknowledgments
Figure 1: Cross Validation performance on FeCrAl Alloy dataset using (a) GP and (b) LVGP models. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainty associated with each prediction
Figure 2: Predicted uncertainty distribution provided by the LVGP and GP-GE models on the 36 sample GE Data testing set
Figure 3: Training and testing performance on FeCrAl Alloy dataset using the LVGP-Target model. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainty associated with each prediction
@@
Figure 5: Predicted uncertainty distribution provided by the LVGP and GP-Prof models on the 6 sample Prop of SmCeCoFeCu testing set